### v1

# Fast ECoT: Efficient Embodied Chain-of-Thought via Thoughts Reuse

Abstract
Embodied Chain-of-Thought (ECoT) reasoning enhances vision-language-action (VLA) models by improving performance and interpretability through intermediate reasoning steps. However, its sequential autoregressive token generation introduces significant inference latency, limiting real-time deployment. We propose Fast ECoT, an inference-time acceleration method that exploits the structured and repetitive nature of ECoT to (1) cache and reuse high-level reasoning across timesteps and (2) parallelise the generation of modular reasoning steps. Additionally, we introduce an asynchronous scheduler that decouples reasoning from action decoding, further boosting responsiveness. Fast ECoT requires no model changes or additional training and integrates easily into existing VLA pipelines. Experiments in both simulation (LIBERO) and real-world robot tasks show up to a 7.5× reduction in latency with comparable or improved task success rate and reasoning faithfulness, bringing ECoT policies closer to practical real-time deployment.
Keywords: Embodied Chain-of-Thought Reasoning, Inference Acceleration, Vision-Language-Action Models
1 Introduction
Large-scale vision-language-action (VLA) models have recently advanced robotic control by leveraging internet-scale visual and textual knowledge [1, 2, 3]. By combining pre-trained vision-language backbones with policy learning, these models exhibit impressive generalisation across open-world tasks. Among their most powerful capabilities is chain-of-thought (CoT) reasoning—the ability to iteratively generate intermediate reasoning steps before taking an action. In robotics, Embodied Chain-of-Thought (ECoT) [4] extends this concept by enabling robots to “think out loud”—generating step-by-step textual reasoning traces (e.g., plans, subgoals, grounded visual inferences) at each time step, explicitly encoding the robot’s thought process before emitting an action. This augmentation improves model interpretability and boosts success rates.
While ECoT offers these benefits, they come at a steep computational cost. Generating reasoning traces involves producing dozens of tokens autoregressively at each time step, resulting in sequential generation delays. As tasks grow more complex, the length of these reasoning chains increases, compounding the latency. In real-time robotic control, policies must react quickly to new observations; the inference overhead introduced by ECoT can slow the control loop to impractical speeds. In other words, the robot idles much of its time “thinking” rather than acting. Reducing this latency without sacrificing reasoning quality or task performance is essential for making ECoT viable in real-world deployments.
In this work, we address this inference bottleneck by proposing Fast ECoT, a novel method for accelerating embodied chain-of-thought reasoning through thought reuse and parallelised reasoning. Our key insight is that ECoT outputs structurally and exhibits a high degree of temporal locality: many reasoning steps—such as recalling the task goal or rechecking the state of a target object—are repeated across time steps. Rather than regenerating the full reasoning trace at every step, we identify and cache recurring reasoning segments, reusing them in subsequent time steps. This reduces redundant computation while preserving the structure and interpretability of the thought process.
Building on this reuse, we introduce a partial parallelisation strategy that transforms the traditionally sequential reasoning process into a batched one. By branching only when necessary to handle novel information, Fast ECoT enables multiple reasoning steps to be generated in parallel—substantially reducing inference latency. We further introduce an asynchronous scheduling mechanism that decouples action and reasoning generation. Recognising that robot actions evolve more slowly than sensor observations, we prioritise fast action decoding while allowing reasoning traces to update asynchronously in the background. This design reduces latency without compromising decision quality, as the cached reasoning remains stable over short time horizons.
In summary, our contributions are as follows:
  • We propose Fast ECoT, a novel inference-time acceleration method that caches and reuses intermediate reasoning steps, enabling partially parallel chain-of-thought generation for the first time in embodied policies.
  • We introduce an asynchronous scheduling strategy that decouples reasoning and action generation, allowing high-frequency control signals to be emitted while reasoning is updated in the background.
  • Our method requires no model changes or additional training and can be seamlessly integrated into existing VLA models that perform embodied reasoning, making it broadly usable in practical robotic systems.
2 Related Work
Foundation Models for Robotic Manipulation. Recent advances in robot learning have produced large-scale generalist policies to excel at robotic manipulation tasks, by first pre-training on diverse multimodal datasets [1, 3, 2], and further fine-tuning on extensive robot-specific data collections [5, 6]. Vision-language-action (VLA) architectures [1, 3, 2]—which integrate vision-language models pretrained on internet-scale corpora [7, 8]—further unify perception, language, and control into a single transformer-based policy, demonstrating state-of-the-art performance on benchmarks. Nonetheless, current VLA approaches purely generate robot actions and largely underutilise the inherent multi-step reasoning capabilities of their language backbones.
Reasoning for Robotic Control. Chain-of-thought (CoT) prompting [9, 10] has proven effective in enhancing large language models by encouraging step-by-step reasoning. Prior methods employ pre-trained language models for high-level planning [11, 12, 13], often requiring separate low-level controllers for execution. More recent work explores end-to-end integration of reasoning into robotic policies. Notably, Embodied Chain-of-Thought (ECoT) [4] enables robots to generate grounded, interpretable reasoning traces referencing visual and environmental observations, leading to improved performance when paired with VLA models like OpenVLA [1]. EMMAX [14] similarly embeds reasoning into action pipelines. Our work builds on ECoT, targeting its major bottleneck—high inference latency from autoregressive token generation—by proposing a partially parallelised reasoning framework that preserves interpretability while substantially accelerating inference.
Inference Optimisation in Language and Multimodal Models. A wide range of techniques have been proposed to speed up inference in autoregressive models. Speculative decoding [15, 16] accelerates generation by predicting tokens with a lightweight draft model and verifying them with the full model. Non-autoregressive and parallel decoding strategies have also been adopted, particularly in machine translation and, more recently, in robotic control [17, 18]. Quantisation methods [19, 20, 21] are widely used to reduce model precision for faster computation. In contrast to these methods, which typically operate at the token or model level, our work focuses on reasoning-level acceleration—breaking CoT generation into reusable, semantically meaningful segments that can be cached and executed in parallel. This approach is model-agnostic and does not rely on auxiliary models, fine-tuning, or precision reduction. By enabling efficient reuse and branching during reasoning, our method offers a lightweight and integrable solution for speeding up VLA policies with CoT reasoning, without compromising interpretability or task success.
3 Preliminaries
Figure 1: ECoT [4] reasoning autoregressively generates high-level (green) and low-level (purple) reasoning steps to enhance VLA performance.
3.1 Embodied Chain-of-Thought Reasoning (ECoT)
Vision‑Language‑Action (VLA) models [1] build on large pre‑trained vision‑language models, fine‑tuning them to map a natural language instruction $I^{t}$ and image observation $O^{t}$ directly to low‑level robot actions $A^{t}$ via autoregressive token prediction at each control step $t$. Embodied Chain‑of‑Thought (ECoT) [4] reasoning augments this reactive paradigm by supervising the model to generate a structured sequence of $N$ intermediate reasoning steps $R^{t}=\{\,r_{i}^{t}\mid i=1,2,\dots,N\}$, e.g. rephrased task, high‑level plan, grounded sub‑task, low‑level move command, visual features, before emitting the final action $A^{t}$. The reasoning steps are separated into high- and low-level steps, as clarified in Fig. 1 and Sec. A.1.
3.2 Continuous Batching
In autoregressive models, batching improves throughput, but static batching can be inefficient when sequence lengths vary—shorter sequences must wait for longer ones, wasting compute on padding. Continuous batching [22] addresses this by dynamically replacing completed sequences with new ones, maintaining high GPU utilization and minimising idle time. This strategy, used in systems like Tensorrt-LLM [23] and vLLM [24], has shown 2–4× throughput gains in LLM serving.
4 Method
(a) Average update ratio of reasoning steps, highlighting that high‑level reasoning components remain relatively stable, whereas lower‑level steps require more frequent updates.
(b) Average length and standard deviation of different reasoning steps, showing significant disparities in token counts.
Figure 2: Statistics illustrating the pattern of ECoT reasoning steps under Bridge V2 [5].
4.1 Inference Characteristics of ECoT Reasoning
Unlike traditional chain-of-thought approaches in language modelling [25, 26] and vision-language modelling [27, 28], which yield diverse, dynamic reasoning patterns that rarely recur, ECoT consistently follows a structured, periodically recurring workflow: planning, sub-task identification, motion reasoning, and visual-feature processing, before predicting subsequent robot actions. To analyse this behaviour quantitatively, we sample all 1,110 episodes containing reasoning from the Bridge V2 dataset [5] and compute the average and the standard deviation of the update ratio (percentage of reasoning content updated at the next time step) and token length for each reasoning step. As depicted in Fig. 2(a), higher-level reasoning components in ECoT (e.g., planning and subtask reasoning) remain relatively similar across multiple time steps within an episode [^1] While profiling the update ratios, we observed substantial variability in the visible detection module. This counterintuitive result stems from the instability inherent in the visual reasoning module of ECoT and the open-vocabulary nature of object labels, whereby identical objects may be assigned varying labels over time.. For example, the planning module exhibits an average update ratio of only $8.4\%$, meaning $91.6\%$ of its reasoning content is unchanged and can be reused in subsequent inference steps without autoregressive regeneration. Leveraging this temporal locality, we propose to cache ECoT reasoning for reuse in successive time steps, thus potentially enabling the parallel generation of each reasoning step.
Figure 3: Comparison between ECoT (left) and our proposed Fast ECoT (right). Both decompose reasoning into fixed stages (e.g., task, plan, object grounding), but ECoT generates these sequentially at every step, while Fast ECoT enables parallel generation and reuses cached high-level reasoning from previous timesteps. The dotted lines coloured in green/magenta represent token copying.
4.2 Efficient Parallelisation of ECoT Reasoning and Action Generation
Compared to the original ECoT—which generates the full reasoning sequence sequentially and autoregressively at every timestep (see Fig. 3 left)—our Fast ECoT reformulates each reasoning step $r_{n}^{t}$ as a standalone generation task. For each step, we construct the input by combining the current observation $O^{t}$, instruction $I^{t}$, and the previously generated reasoning steps $R^{t}=\{r_{i}^{t}\mid i=1,2,\dots,n-1\}$ from the last timestep as prefix context. This allows all reasoning steps and the action at timestep $t$ to be generated independently and in parallel, rather than waiting for preceding components to finish (see Fig. 3 right).
While conceptually straightforward, this strategy introduces performance overhead. Since each reasoning module prepends a growing context of prior thoughts, the resulting input lengths vary significantly—early steps have short prompts, while later ones accumulate more history. Additionally, output lengths also differ across reasoning steps: low-level steps like gripper commands may require fewer than 20 tokens, while object-grounding reasoning can exceed 120 (see Fig. 2(b)). Traditional static batching [22] (see Fig. 4) handles such variability by padding all sequences in a batch to match the longest, which leads to substantial inefficiency on modern accelerators. This results in wasted compute on padding tokens, poor GPU utilisation, and offsets the gains from parallel generation.
To address inefficiencies from padding, we adopt continuous batching [22], a dynamic scheduling strategy used in modern LLM serving engines[24, 23]. Instead of padding all sequences to a fixed length, continuous batching maintains a queue where completed sequences are immediately replaced by new ones, allowing variable-length inputs to be processed efficiently. This minimises wasted computation on padding tokens, significantly improves GPU utilisation, and can reduce the total number of tokens processed by up to 6× (see Fig. 4). We use vLLM [24] as our inference backend, and the full pseudocode for Fast ECoT is provided in Sec. A.2.
Figure 4: Illustrative example comparing static vs. continuous batching for reasoning generation. Left: Static batching pads to the longest sequence, processing 4×11=44 tokens. Right: Continuous batching processes only actual tokens (3, 6, 8, 9), adding up to 26 tokens, reducing padding and improving efficiency.
4.3 Optimising Asynchronous Scheduling of Reasoning and Action Updates
Figure 5: Generated robot rollouts at successive time steps (top row) alongside its reasoning (bottom row). Between frames, a large part of the reasoning remains unchanged (Green). At each timestep (t=1, 10, 24, 31, 33, 42), the Subtask updates intermittently, and the low-level Move command adapts continuously as it picks up the banana and places it on the plate.
As discussed in Sec. 4.2, we formulate reasoning and action generation at each time step as batched requests that can be processed in parallel. However, reasoning traces typically span hundreds of tokens, while action decoding involves only a few (around $7$) tokens. In a synchronised setup, this mismatch causes unnecessary latency: the agent must wait for all reasoning steps to complete before it can act.
While the original ECoT framework couples reasoning and action tightly for interpretability and causal grounding, we argue that this coupling need not be time-synchronous. In practice, high-level reasoning elements—such as Task and Plan—remain stable across multiple steps, and their semantic influence on action decisions persists even when updated less frequently. This is evident in the rollout shown in Fig. 5, where reasoning traces and visible objects stay largely consistent despite continuously evolving actions. Motivated by this temporal stability, we adopt an asynchronous design that reuses cached high-level reasoning across timesteps, avoiding unnecessary regeneration. This is particularly beneficial in robotic settings, where action frequencies are faster than reasoning updates. By decoupling reasoning from action generation and updating the cached reasoning trace $R^{c}$ in the background, our approach enables low-latency action decoding without compromising interpretability or task performance. The asynchronous scheduling mechanism separately handles reasoning and action requests, returning actions immediately upon decoding, while reasoning is refreshed opportunistically. Full pseudocode is provided in Alg. 1.
Algorithm 1 Asynchronous Parallel Embodied Chain‑of‑Thought
1:Time step $t$, Instruction $I^{t}$, Observation $O^{t}$, History Reasoning $R^{c}$
2:$c^{t}\leftarrow\mathrm{encode}(I^{t},O^{t})$
3:for $i=1\,\mathbf{to}\,N+1$ concurrently do
4:     Lock $R^{c}$ and autoregressively sample $r_{i}^{t}\sim P\bigl{(}r_{i}^{t}\mid c^{t},R^{c}[:i]\bigr{)}$
5:     if $i=N+1$ then
6:         $A^{t}=r_{i}^{t}$
7:     else
8:         Lock $R^{c}$ and update $R^{c}$ with $r_{i}^{t}$
9:     end if
10:end for
11:wait until $A^{t}$ is finished
12:$A^{t}\leftarrow\mathrm{decode}(A^{t})$
13:return $R^{c},A^{t}$
5 Experimental Results
In this section, we conduct experiments to evaluate the effectiveness of Fast ECoT. Our evaluation focuses on the following key questions: (1) To what extent does parallel reasoning generation improve computational efficiency? (2) Does the proposed method preserve task performance comparable to the sequential baseline? (3) What is the impact of reasoning step parallelization on the overall quality of reasoning?
5.1 LIBERO Simulation Experimental Setup
Task Setup. We conduct experiments using a Franka Emika Panda robotic arm within the LIBERO environment [29], a widely adopted benchmark for evaluating generalizable robotic policies. To comprehensively assess policy generalisation, we select four diverse task suites—LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long—each targeting distinct challenges, including spatial configuration, object manipulation, goal specification, and long-horizon task execution.
Training Data and Training Recipe. ECoT models require fine-tuning on each task for optimal performance [1, 4]. Training data is generated following the ECoT pipeline [4], with modifications for reproduction: we integrate Janus [30] for automated scene description and Deepseek-Reasoner [26] for generating CoT reasoning trajectories from successful demonstrations. We initialise ECoT models from the open-sourced checkpoint [4], which was pre-trained on Bridge V2 [5] and OXE dataset [6]. Then we apply LoRA [31] with rank 32 and train for $200,000$ gradient steps using a batch size of 1 distributed across 4 NVIDIA A6000 GPUs. See Sec. B.1 for other hyperparameters.
Baselines. We compare our proposed method, Fast ECoT, against four baselines: ECoT, the original ECoT model that autoregressively generates the full reasoning chain; ECoT (5-step), a variant that updates low-level reasoning at every timestep but updates high-level reasoning only every 5 timesteps; ECoT (async), a variant originally designed to use two GPUs to asynchronously compute high-level and low-level reasoning, which we adapt to run entirely on a single GPU for fair comparison; and ECoT (Quant), a post-training quantized version of ECoT utilizing Huggingface’s Bitsandbytes [32], selected due to its best acceleration performance compared to other quantization approaches tested [19, 20]. For fairness, we exclude methods requiring additional training or architectural changes, such as speculative decoding [15, 16], and action chunking [18, 33].
5.2 LIBERO Simulation Experimental Results
Method LIBERO-Object SR (%) $\uparrow$ LIBERO-Spatial SR (%) $\uparrow$ LIBERO-Goal SR (%) $\uparrow$ LIBERO-Long SR (%) $\uparrow$ Average SR (%) $\uparrow$ Latency Per Step (ms) $\downarrow$
ECoT 77 84 75 45 70.3 4997 $\pm$ 691
ECoT (5-step) 79 75 72 40 66.5 3514 $\pm$ 969
ECoT (Async) 70 83 80 51 71.0 3655 $\pm$ 773
ECoT (Quant) 82 82 84 53 75.3 2180 $\pm$ 207
Fast ECoT 83 85 83 69 80.0 2156 $\pm$ 353
Fast ECoT (Async) 75 83 83 69 77.5 686 $\pm$ 412
Table 1: LIBERO simulation experimental results. SR = Success Rate. All results are averaged over 100 trials per suite and run on a single NVIDIA RTX 4090 GPU. Best results are in bold.
Inference Speedup. As shown in the last column of Tab. 1, our parallel reasoning strategy significantly accelerates inference latency. Specifically, Fast ECoT achieves a latency reduction to $2156\pm 353$ ms per step, representing over a 2.3x speed-up compared to the original ECoT baseline ($4997\pm 691$ ms). The asynchronous variant, Fast ECoT (Async), further reduces latency to $686\pm 412$ ms, nearly a 7x improvement. Notably, both variants maintain relatively stable latency compared to baselines proposed in ECoT [4].
Task Success Rate. Despite the significant reduction in latency, our method preserves or enhances task performance relative to baseline methods (see Tab. 1). Fast ECoT achieves the highest overall performance, with an average success rate (SR) of $80.0\%$, exceeding the baseline ECoT in all LIBERO suites. We hypothesise that our parallel reasoning strategy—which incorporates prior reasoning steps—helps smooth temporal inconsistencies and contributes to this performance gain. Notably, Fast ECoT (Async) also achieves a comparable overall performance, with an average SR of $77.5\%$, including the best result on LIBERO-Long ($69\%$). The largest decline compared with normal Fast ECoT happens at LIBERO-Object (from 83% to 75%), where the testing scenes focus on diverse layouts of the object. The positions of visible objects lead to a greater impact on overall control performance, and the larger temporal mismatch between reasoning components and actions of Fast ECoT (Async) may cause this degradation. Surprisingly, ECoT (Quant) also shows decent performance improvements compared to the original ECoT, which can be understood as regularising model behaviour in testing environments [34].
5.3 Real-world Experiments
Setup. We validate Fast ECoT on a physical Franka Emika Panda robotic arm equipped with a RealSense D455 camera, providing third-person RGB-D observations. We design six manipulation tasks representative of common household scenarios. The evaluation includes both in-distribution and out-of-distribution tasks featuring unseen objects and instructions, detailed in Sec. B.2. We collect 50 expert demonstrations via Droid teleoperation pipeline [35] to fine-tune ECoT models. We follow the same training data generation process, training recipe, and baselines used in Sec. 5.1.
Method ID SR (%) $\uparrow$ OOD Objects SR (%) $\uparrow$ OOD Instruction SR (%) $\uparrow$ Average SR (%) $\uparrow$ Latency per Step (ms) $\downarrow$
ECoT 78.3 73.3 40.0 64.0 5556 $\pm$ 384
ECoT (5-step) 56.6 51.7 35.0 47.8 4327 $\pm$ 619
ECoT (Async) 76.6 70.0 41.7 62.8 4206 $\pm$ 323
ECoT (Quant) 75.0 63.3 48.3 62.2 2437 $\pm$ 171
Fast ECoT 76.6 66.7 46.7 63.3 2479 $\pm$ 520
Fast ECoT (Async) 83.3 76.7 50.0 70.0 716 $\pm$ 529
Table 2: Real-world experimental results on selected household tasks. SR = Success Rate. ID = In distribution. OOD = Out of distribution. All results are averaged over 180 trials and run on a single NVIDIA RTX 4090 GPU. Best results are in bold.
Results. Tab. 2 summarises the performance of our method against baseline approaches on real-world tasks. Fast ECoT and Fast ECoT (Async) both demonstrate substantial latency reductions while preserving or even improving task success rates. Specifically, Fast ECoT (Async) delivers the best performance, with an average success rate of 70% across in-distribution (ID), out-of-distribution (OOD) objects, and OOD instruction tasks, alongside the lowest latency of $716\pm 529$ ms per step—a 7.5x speedup over the original ECoT baseline ($5556\pm 384$ ms). This result reinforces the benefits of asynchronous parallel reasoning in real-world deployment scenarios, where low-variance and rapid inference are critical.
5.4 Action Faithfulness in ECoT Reasoning
Figure 6: Action faithfulness (AF) on LIBERO tasks. All graphs are plotted with mean and standard deviation shading across 1000 steps.
To validate that the generated chain-of-thought (CoT) not only improves performance but also faithfully explains the model’s decision-making process, we build on the faithfulness criteria for language models [36] and introduce a novel quantitative metric - action faithfulness (AF) - to measure the faithfulness of the CoT reasoning for robotic tasks. fFormally, given a complete CoT consisting of $N$ intermediate reasoning steps that culminate in a final robot action $A$, we enforce the model to directly predict an action $A_{i}$ after generating only the first $i$ reasoning steps, where $i=0,1,\dots,N$. We then compute the L1 distance between $A_{i}$ and $A$ as the faithfulness score $AF_{i}=\|A_{i}-A\|_{1}$. Higher L1 distances indicate a greater dependence on subsequent reasoning steps, thereby suggesting higher CoT faithfulness.
We plot action faithfulness scores of Fast ECoT and ECoT for all reasoning steps $\{AF_{i}\}$ in Fig. 6. Fast ECoT preserves the faithfulness of the base ECoT model during parallel reasoning generation for most reasoning steps. The action faithfulness without predicting “Visible Objects” is slightly lower for Fast ECoT, since the high update ratio and one-step delay in visual features might increase its chances to be post-hoc. Meanwhile, the asynchronous variant of Fast ECoT generally exhibits a lower faithfulness, probably due to the larger temporal mismatch introduced in async reasoning. Faithfulness is notably low when no reasoning steps are generated (i.e., at “Task”), likely due to fine-tuning from the OpenVLA [1] checkpoint, which did not require reasoning steps. We provide more CoT reasoning examples in Sec. C.1.
6 Conclusion
We present Fast ECoT, an inference-time acceleration method for Embodied Chain-of-Thought (ECoT) reasoning in VLA models. By exploiting structural and temporal locality, Fast ECoT enables (1) reuse of cached high-level reasoning and (2) parallel generation of modular steps. An asynchronous scheduler further decouples reasoning from action decoding, enhancing real-time responsiveness. Requiring no architectural changes or retraining, Fast ECoT integrates easily into existing VLA pipelines. Across simulations and real-world tests, it reduces inference latency by up to 7.5× while maintaining task performance and interpretability, advancing ECoT toward real-time deployment.
7 Limitations
While our work demonstrated that ECoT reasoning can be significantly accelerated through the reuse of previously computed thoughts, there remain several limitations. First, we still update all reasoning steps at once without identifying and prioritising the more vital steps for frequent updates. This approach can lead to excessive computation, and more sophisticated methods could dynamically determine which reasoning steps require updating based on their relative importance to task success. Second, the effectiveness of thoughts reuse inherently relies on the assumption that at least part of the reasoning remains stable over a short time window. In scenarios where reasoning steps frequently change between consecutive time steps, this stability assumption breaks down, potentially degrading the performance gains offered by our method. Third, our current approach assumes a fixed structure in reasoning chains, thus limiting its applicability to tasks or environments requiring highly dynamic or adaptive reasoning patterns. Extending thoughts reuse to accommodate more diverse and adaptive reasoning structures remains an open challenge. Addressing these limitations represents promising avenues for future research to further enhance the efficiency and generality of embodied reasoning in robotic control.
Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
Appendix A Extra Algorithm Details
A.1 Algorithm on Embodied Chain-of-Thought
Algorithm 2 Embodied Chain‑of‑Thought
1:Time step $t$, Instruction $I^{t}$, Observation $O^{t}$
2:$c^{t}\leftarrow\mathrm{encode}(I^{t},O^{t})$
3:$R^{t}\leftarrow[\,]$
4:for $i=1\,\mathbf{to}\,N+1$ do
5:     Autoregressively sample $r_{i}^{t}\sim P\bigl{(}r_{i}^{t}\mid c^{t},R^{t}\bigr{)}$
6:     Add $r_{i}^{t}$ to $R^{t}$
7:end for
[…]
9:return $R^{t},A^{t}$
A.2 Algorithm on Parallel Embodied Chain-of-Thought
Algorithm 3 Parallel Embodied Chain‑of‑Thought
1:Time step $t$, Instruction $I^{t}$, Observation $O^{t}$, Last reasoning steps $R^{t-1}$
2:$c^{t}\leftarrow\mathrm{encode}(I^{t},O^{t})$
3:$R^{t}\leftarrow[\,]$
4:for $i=1\,\mathbf{to}\,N+1$ concurrently do
5:     Autoregressively sample $r_{i}^{t}\sim P\bigl{(}r_{i}^{t}\mid c^{t},R^{t-1}[:i]\bigr{)}$
6:end for
7:Synchronize
[…]
9:return $R^{t},A^{t}$
Appendix B Extra Experimental Setup
B.1 Finetuning Hyperparameters
Hyperparameter Value
Base Model ecot-openvla-7b-oxe
Number of GPUs 4
Batch Size 1
Gradient Accumulation Steps 1
Learning Rate 5e-4
Learning Rate Scheduler constant LR
Optimizer AdamW
AdamW Betas (0.9, 0.999)
AdamW Epsilon 1e-8
Training Steps 200,000
Image Augmentation
Reasoning Dropout 0.0
LoRA Rank 32
LoRA Alpha 16
LoRA Dropout 0.0
Table 3: Fine-tuning hyperparameters used for fine-tuning EcoT model.
B.2 Real-World Experimental Setup
Out-of-distribution Instructions. We use GPT-4o [37] to rephrase instructions from the training set. These rephrasings modify both the verbs and nouns to create linguistically novel yet semantically equivalent instructions, allowing us to test generalization to unseen language.
Original Instruction OOD Rephrased Instruction
Pick up the gray spanner and put it in the basket Grab the silver wrench and drop it into the bin
Pick up the blue hammer and put it in the basket Take the cobalt mallet and place it in the container
Pick up the blue squash ball and put it in the basket Retrieve the navy rubber ball and set it in the basket
Pick up the white baseball and put it in the basket Grab the pale softball and store it in the bucket
Pick up the banana and put it on the plate Take the yellow fruit and lay it on the dish
Pick up the lemon and put it in the cup Seize the citrus and drop it into the mug
Table 4: Original instructions and their corresponding out-of-distribution (OOD) rephrasings generated by GPT-4o.
In-distribution Objects.
Figure 7: Qualitative examples of in-distribution (ID) objects used for evaluation.
Out-of-distribution objects.
Figure 8: Qualitative examples of out-of-distribution (OOD) objects used for evaluation.
Appendix C Extra Experimental Results
C.1 CoT Reasoning Examples
Figure 9: Qualitative examples of ECOT reasoning in real-world experiments.
Figure 10: Qualitative examples of ECOT reasoning in simulation. As the VLMs used in both the data generation pipeline and the underlying VLA model are pretrained on real-world images, their performance in simulated environments is less stable. This can lead to inaccurate bounding boxes, undetected objects, and unusual or incorrect object labels.

### v2

@@
 Additionally, we introduce an asynchronous scheduler that decouples reasoning from action decoding, further boosting responsiveness.
-Fast ECoT requires no model changes or additional training and integrates easily into existing VLA pipelines.
+Fast ECoT requires no model changes or additional training and easily integrates into existing VLA pipelines.
 Experiments in both simulation (LIBERO) and real-world robot tasks show up to a 7.5× reduction in latency with comparable or improved task success rate and reasoning faithfulness, bringing ECoT policies closer to practical real-time deployment.
-Keywords: Embodied Chain-of-Thought Reasoning, Inference Acceleration, Vision-Language-Action Models
-1 Introduction
+The code will be released upon acceptance.
+I INTRODUCTION
 Large-scale vision-language-action (VLA) models have recently advanced robotic control by leveraging internet-scale visual and textual knowledge [1, 2, 3].
@@
 We further introduce an asynchronous scheduling mechanism that decouples action and reasoning generation.
-Recognising that robot actions evolve more slowly than sensor observations, we prioritise fast action decoding while allowing reasoning traces to update asynchronously in the background.
+Recognising that robot actions evolve faster than reasoning traces, we prioritise fast action decoding while allowing reasoning traces to update asynchronously in the background.
 This design reduces latency without compromising decision quality, as the cached reasoning remains stable over short time horizons.
-In summary, our contributions are as follows:
-  • We propose Fast ECoT, a novel inference-time acceleration method that caches and reuses intermediate reasoning steps, enabling partially parallel chain-of-thought generation for the first time in embodied policies.
-  • We introduce an asynchronous scheduling strategy that decouples reasoning and action generation, allowing high-frequency control signals to be emitted while reasoning is updated in the background.
-  • Our method requires no model changes or additional training and can be seamlessly integrated into existing VLA models that perform embodied reasoning, making it broadly usable in practical robotic systems.
-2 Related Work
+II RELATED WORK
+Figure 1: ECoT [4] reasoning autoregressively generates high-level (green) and low-level (purple) reasoning steps to enhance VLA performance.
 Foundation Models for Robotic Manipulation.
-Recent advances in robot learning have produced large-scale generalist policies to excel at robotic manipulation tasks, by first pre-training on diverse multimodal datasets [1, 3, 2], and further fine-tuning on extensive robot-specific data collections [5, 6].
-Vision-language-action (VLA) architectures [1, 3, 2]—which integrate vision-language models pretrained on internet-scale corpora [7, 8]—further unify perception, language, and control into a single transformer-based policy, demonstrating state-of-the-art performance on benchmarks.
-Nonetheless, current VLA approaches purely generate robot actions and largely underutilise the inherent multi-step reasoning capabilities of their language backbones.
+Recent advances in robot learning have produced large-scale generalist policies that excel at robotic manipulation tasks, by first pre-training on diverse multimodal datasets [1, 2, 3], and further fine-tuning on extensive robot-specific data collections [5, 6].
+Vision-language-action (VLA) architectures [1, 2, 3]—which integrate vision-language models pretrained on internet-scale corpora [7, 8]—unify perception, language, and control into a single transformer-based policy, achieving state-of-the-art performance.
+Current VLA methods can be broadly categorized into two families: (1) monolithic models, which integrate perception, language, and action within either single- or dual-system architectures [1, 2, 3]; and (2) hierarchical models, which explicitly decouple planning from policy execution by producing interpretable intermediate representations such as subtasks, keypoints, or programs [9, 10].
+While monolithic approaches highlight simplicity and end-to-end generalisation, hierarchical designs emphasise interpretability and modularity for long-horizon control.
 Reasoning for Robotic Control.
@@
 Prior methods employ pre-trained language models for high-level planning [13, 14, 15], often requiring separate low-level controllers for execution.
-More recent work explores end-to-end integration of reasoning into robotic policies.
-Notably, Embodied Chain-of-Thought (ECoT) [4] enables robots to generate grounded, interpretable reasoning traces referencing visual and environmental observations, leading to improved performance when paired with VLA models like OpenVLA [1].
-EMMAX [14] similarly embeds reasoning into action pipelines.
-Our work builds on ECoT, targeting its major bottleneck—high inference latency from autoregressive token generation—by proposing a partially parallelised reasoning framework that preserves interpretability while substantially accelerating inference.
+Recent work begins to integrate explicit reasoning end-to-end: ECoT introduces embodied chain-of-thought traces grounded in observations [4], CoT-VLA predicts visual subgoal observations [16], and RAD [17] and ThinkAct [18] curate or align language reasoning for low-level action; however, these often rely on latent embeddings, generated sub-goals, or textual descriptions that are hard to precisely ground for manipulation.
+EMMAX [19] embeds reasoning as subtasks and predicted gripper states (2D/3D), but leverages limited full-scene context.
+In contrast, MolmoAct [20] performs “reasoning in space,” explicitly grounding each step in the scene so it can be decoded and visualised on the image plane and within the 3D environment, improving explainability and action prediction.
+Orthogonal to reasoning design, efficiency-oriented VLAs such as Spec-VLA [21] and FlashVLA [22] boost responsiveness by replacing strictly sequential decoding with speculative, parallel, or retraining-free acceleration.
+Our work builds on the ECoT but targets its chief bottleneck—autoregressive latency—via a partially parallelised reasoning framework that preserves interpretability while substantially reducing inference time.
 Inference Optimisation in Language and Multimodal Models.
-A wide range of techniques have been proposed to speed up inference in autoregressive models.
+A wide range of techniques has been proposed to speed up inference in autoregressive models.
 Speculative decoding [23, 24] accelerates generation by predicting tokens with a lightweight draft model and verifying them with the full model.
@@
 By enabling efficient reuse and branching during reasoning, our method offers a lightweight and integrable solution for speeding up VLA policies with CoT reasoning, without compromising interpretability or task success.
-3 Preliminaries
-Figure 1: ECoT [4] reasoning autoregressively generates high-level (green) and low-level (purple) reasoning steps to enhance VLA performance.
-3.1 Embodied Chain-of-Thought Reasoning (ECoT)
+III PRELIMINARIES
+III-A Embodied Chain-of-Thought Reasoning (ECoT)
 Vision‑Language‑Action (VLA) models [1] build on large pre‑trained vision‑language models, fine‑tuning them to map a natural language instruction $I^{t}$ and image observation $O^{t}$ directly to low‑level robot actions $A^{t}$ via autoregressive token prediction at each control step $t$.
@@
 rephrased task, high‑level plan, grounded sub‑task, low‑level move command, visual features, before emitting the final action $A^{t}$.
-The reasoning steps are separated into high- and low-level steps, as clarified in Fig.
-1 and Sec.
-A.1.
-3.2 Continuous Batching
+The reasoning steps are separated into high- and low-level steps (see Fig.
+1).
+III-B Continuous Batching
 In autoregressive models, batching improves throughput, but static batching can be inefficient when sequence lengths vary—shorter sequences must wait for longer ones, wasting compute on padding.
 Continuous batching [29] addresses this by dynamically replacing completed sequences with new ones, maintaining high GPU utilization and minimising idle time.
-This strategy, used in systems like Tensorrt-LLM [23] and vLLM [24], has shown 2–4× throughput gains in LLM serving.
-4 Method
-(a) Average update ratio of reasoning steps, highlighting that high‑level reasoning components remain relatively stable, whereas lower‑level steps require more frequent updates.
-(b) Average length and standard deviation of different reasoning steps, showing significant disparities in token counts.
-Figure 2: Statistics illustrating the pattern of ECoT reasoning steps under Bridge V2 [5].
-4.1 Inference Characteristics of ECoT Reasoning
+This strategy, adopted by the work [30, 31], has shown 2–4× throughput gains in LLM serving.
+IV Method
+(a) Mean update ratio ($\pm$1 standard deviation) per reasoning step, showing high-level stability versus frequent low-level updates.
+(b) Mean length ($\pm$1 standard deviation) of different reasoning steps, showing significant disparities in token counts.
+Figure 2: Statistics illustrating the pattern of ECoT reasoning steps under Libero Goal [32] and Bridge V2 [5].
+IV-A Inference Characteristics of ECoT Reasoning
 Unlike traditional chain-of-thought approaches in language modelling [33, 34] and vision-language modelling [35, 36], which yield diverse, dynamic reasoning patterns that rarely recur, ECoT consistently follows a structured, periodically recurring workflow: planning, sub-task identification, motion reasoning, and visual-feature processing, before predicting subsequent robot actions.
-To analyse this behaviour quantitatively, we sample all 1,110 episodes containing reasoning from the Bridge V2 dataset [5] and compute the average and the standard deviation of the update ratio (percentage of reasoning content updated at the next time step) and token length for each reasoning step.
+To analyse this behaviour quantitatively, we sample all episodes containing reasoning from the Bridge V2 dataset [5] and compute the average and the standard deviation of the update ratio (percentage of reasoning content updated at the next time step) and token length for each reasoning step.
 As depicted in Fig.
@@
 Leveraging this temporal locality, we propose to cache ECoT reasoning for reuse in successive time steps, thus potentially enabling the parallel generation of each reasoning step.
+We observe a similar trend in the simulated environment, LIBERO-Goal [32]: high-level reasoning updates rarely while low-level components refresh more frequently (see Fig.
+2); This consistency supports caching high-level reasoning and densely updating low-level content.
 Figure 3: Comparison between ECoT (left) and our proposed Fast ECoT (right).
-Both decompose reasoning into fixed stages (e.g., task, plan, object grounding), but ECoT generates these sequentially at every step, while Fast ECoT enables parallel generation and reuses cached high-level reasoning from previous timesteps.
+Both decompose reasoning into fixed stages (e.g., task, plan, object grounding), but ECoT generates these sequentially at every step, while Fast ECoT enables parallel generation and reuses cached higher-level reasoning from previous timesteps as context.
 The dotted lines coloured in green/magenta represent token copying.
-4.2 Efficient Parallelisation of ECoT Reasoning and Action Generation
+IV-B Parallelising Reasoning and Action Generation
 Compared to the original ECoT—which generates the full reasoning sequence sequentially and autoregressively at every timestep (see Fig.
 3 left)—our Fast ECoT reformulates each reasoning step $r_{n}^{t}$ as a standalone generation task.
-For each step, we construct the input by combining the current observation $O^{t}$, instruction $I^{t}$, and the previously generated reasoning steps $R^{t}=\{r_{i}^{t}\mid i=1,2,\dots,n-1\}$ from the last timestep as prefix context.
+For each step, we construct the input by combining the current observation $O^{t}$, instruction $I^{t}$, and the previously generated reasoning steps $R^{t-1}=\{r_{i}^{t-1}\mid i=1,2,\dots,n-1\}$ from the last timestep as prefix context.
 This allows all reasoning steps and the action at timestep $t$ to be generated independently and in parallel, rather than waiting for preceding components to finish (see Fig.
@@
 4).
-We use vLLM [24] as our inference backend, and the full pseudocode for Fast ECoT is provided in Sec.
-A.2.
-Figure 4: Illustrative example comparing static vs.
+We use vLLM [31] as our inference backend, and the pseudocode of Fast ECoT is shown in Alg.
+1.
+Algorithm 1 Parallel Embodied Chain‑of‑Thought
+1:Time step $t$, Instruction $I^{t}$, Observation $O^{t}$, Last reasoning steps $R^{t-1}$
+[…]
+[…]
+4:for $i=1\,\mathbf{to}\,N+1$ concurrently do
+5:  Autoregressively sample $r_{i}^{t}\sim P\bigl(r_{i}^{t}\mid c^{t},R^{t-1}[:i]\bigr)$
+6:end for
+7:Synchronize
+[…]
+9:return $R^{t},A^{t}$
+Figure 4: Illustratively comparing static vs.
 continuous batching for reasoning generation.
 Left: Static batching pads to the longest sequence, processing 4×11=44 tokens.
-Right: Continuous batching processes only actual tokens (3, 6, 8, 9), adding up to 26 tokens, reducing padding and improving efficiency.
-4.3 Optimising Asynchronous Scheduling of Reasoning and Action Updates
+Right: Continuous batching processes only actual tokens (3, 6, 8, 9), adding up to 26 tokens, which reduces padding and improves efficiency.
+IV-C Asynchronous Reasoning and Action Updates
+So far we formulate reasoning and action generation at each time step as batched requests that can be processed in parallel.
+However, reasoning traces typically span hundreds of tokens, while action decoding involves only a few (around $7$) tokens.
+In a synchronised setup, this mismatch causes unnecessary latency: the agent must wait for all reasoning steps to complete before it can act (see Fig.
+6).
 Figure 5: Generated robot rollouts at successive time steps (top row) alongside its reasoning (bottom row).
@@
 At each timestep (t=1, 10, 24, 31, 33, 42), the Subtask updates intermittently, and the low-level Move command adapts continuously as it picks up the banana and places it on the plate.
-As discussed in Sec.
-4.2, we formulate reasoning and action generation at each time step as batched requests that can be processed in parallel.
-However, reasoning traces typically span hundreds of tokens, while action decoding involves only a few (around $7$) tokens.
-In a synchronised setup, this mismatch causes unnecessary latency: the agent must wait for all reasoning steps to complete before it can act.
-While the original ECoT framework couples reasoning and action tightly for interpretability and causal grounding, we argue that this coupling need not be time-synchronous.
-In practice, high-level reasoning elements—such as Task and Plan—remain stable across multiple steps, and their semantic influence on action decisions persists even when updated less frequently.
-This is evident in the rollout shown in Fig.
-5, where reasoning traces and visible objects stay largely consistent despite continuously evolving actions.
-Motivated by this temporal stability, we adopt an asynchronous design that reuses cached high-level reasoning across timesteps, avoiding unnecessary regeneration.
-This is particularly beneficial in robotic settings, where action frequencies are faster than reasoning updates.
-By decoupling reasoning from action generation and updating the cached reasoning trace $R^{c}$ in the background, our approach enables low-latency action decoding without compromising interpretability or task performance.
-The asynchronous scheduling mechanism separately handles reasoning and action requests, returning actions immediately upon decoding, while reasoning is refreshed opportunistically.
-Full pseudocode is provided in Alg.
-1.
+(a) Synchronous inference.
+(b) Asynchronous inference.
+Figure 6: Illustration of inference stacks.
+Asynchronous overlap reduces stall time (green dashed) of robot operation and increases action throughput within the same time window compared to synchronous execution.
+While ECoT couples reasoning and action for interpretability and causal grounding, this coupling need not be time-synchronous.
+Empirically, high-level elements (e.g., Task, Plan) change slowly across steps; their influence on actions persists even with infrequent updates—see rollouts in Fig.
+5, where reasoning traces and visible objects remain stable as actions evolve.
+Inspired by VLA systems that separate a fast action module from a slower VLM (SmolVLA [37]; GR00T [38]), we adopt an asynchronous split: the controller decodes actions from the current observation $O^{t}$ and a cached high-level reasoning $R^{c}$, while $R^{c}$ is refreshed in the background.
+As shown in Fig.
+6, asynchronous reasoning action generation effectively reduces stalls and increases action throughput without sacrificing interpretability.
+Full pseudocode of the above is shown in Alg.
+2.
 Algorithm 2 Asynchronous Parallel Embodied Chain‑of‑Thought
 1:Time step $t$, Instruction $I^{t}$, Observation $O^{t}$, History Reasoning $R^{c}$
-2:$c^{t}\leftarrow\mathrm{encode}(I^{t},O^{t})$
+[…]
 3:for $i=1\,\mathbf{to}\,N+1$ concurrently do
@@
 11:wait until $A^{t}$ is finished
-12:$A^{t}\leftarrow\mathrm{decode}(A^{t})$
+[…]
 13:return $R^{c},A^{t}$
-5 Experimental Results
+V RESULTS
 In this section, we conduct experiments to evaluate the effectiveness of Fast ECoT.
@@
 (3) What is the impact of reasoning step parallelization on the overall quality of reasoning?
-5.1 LIBERO Simulation Experimental Setup
+V-A LIBERO Experiments
 Task Setup.
@@
 Then we apply LoRA [40] with rank 32 and train for $200,000$ gradient steps using a batch size of 1 distributed across 4 NVIDIA A6000 GPUs.
-See Sec.
-B.1 for other hyperparameters.
 Baselines.
@@
 For fairness, we exclude methods requiring additional training or architectural changes, such as speculative decoding [23, 24], and action chunking [25, 42].
-5.2 LIBERO Simulation Experimental Results
 Method LIBERO-Object SR (%) $\uparrow$ LIBERO-Spatial SR (%) $\uparrow$ LIBERO-Goal SR (%) $\uparrow$ LIBERO-Long SR (%) $\uparrow$ Average SR (%) $\uparrow$ Latency per Step (ms) $\downarrow$
-ECoT 77 84 75 45 70.3 4997 $\pm$ 691
-ECoT (5-step) 79 75 72 40 66.5 3514 $\pm$ 969
-ECoT (Async) 70 83 80 51 71.0 3655 $\pm$ 773
-ECoT (Quant) 82 82 84 53 75.3 2180 $\pm$ 207
+OpenVLA 87 83 74 55 75.3 184 $\pm$ 37
+ECoT 77 84 75 57 73.3 4997 $\pm$ 691
+ECoT (5-step) 79 75 72 56 70.5 3514 $\pm$ 969
+ECoT (Async) 70 83 80 47 70.0 3655 $\pm$ 773
+ECoT (Quant) 82 82 84 57 76.3 2180 $\pm$ 207
 Fast ECoT 83 85 83 69 80.0 2156 $\pm$ 353
 Fast ECoT (Async) 75 83 83 69 77.5 686 $\pm$ 412
-Table 1: LIBERO simulation experimental results. SR = Success Rate. All results are averaged over 100 trials per suite and run on a single NVIDIA RTX 4090 GPU. Best results are in bold.
-Inference Speedup.
-As shown in the last column of Tab.
-1, our parallel reasoning strategy significantly accelerates inference latency.
-Specifically, Fast ECoT achieves a latency reduction to $2156\pm 353$ ms per step, representing over a 2.3x speed-up compared to the original ECoT baseline ($4997\pm 691$ ms).
-The asynchronous variant, Fast ECoT (Async), further reduces latency to $686\pm 412$ ms, nearly a 7x improvement.
-Notably, both variants maintain relatively stable latency compared to baselines proposed in ECoT [4].
-Task Success Rate.
-Despite the significant reduction in latency, our method preserves or enhances task performance relative to baseline methods (see Tab.
-1).
-Fast ECoT achieves the highest overall performance, with an average success rate (SR) of $80.0\%$, exceeding the baseline ECoT in all LIBERO suites.
-We hypothesise that our parallel reasoning strategy—which incorporates prior reasoning steps—helps smooth temporal inconsistencies and contributes to this performance gain.
-Notably, Fast ECoT (Async) also achieves a comparable overall performance, with an average SR of $77.5\%$, including the best result on LIBERO-Long ($69\%$).
-The largest decline compared with normal Fast ECoT happens at LIBERO-Object (from 83% to 75%), where the testing scenes focus on diverse layouts of the object.
-The positions of visible objects lead to a greater impact on overall control performance, and the larger temporal mismatch between reasoning components and actions of Fast ECoT (Async) may cause this degradation.
-Surprisingly, ECoT (Quant) also shows decent performance improvements compared to the original ECoT, which can be understood as regularising model behaviour in testing environments [34].
-5.3 Real-world Experiments
+TABLE I: LIBERO simulation experimental results. SR = Success Rate.
+[…]
+Performance.
+Our method consistently improves task performance relative to prior baselines.
+Fast ECoT achieves the highest average success rate ($80.0\%$), surpassing all ECoT variants and non-reasoning OpenVLA.
+We attribute this to our parallel reasoning strategy, which smooths temporal inconsistencies and leverages prior reasoning steps.
+Fast ECoT (Async) remains competitive ($77.5\%$) and achieves the best result on LIBERO-Long ($69\%$), though at some cost in LIBERO-Object due to greater sensitivity to object layouts and temporal mismatch.
+Interestingly, ECoT (Quant) also improves over vanilla ECoT, likely due to quantisation regularising model behaviour [43].
+Fig.
+7 shows that both ECoT and Fast ECoT (Async) produce coherent, grounded reasoning, but the latter reaches comparable states much quicker.
+(a) ECoT at $t=143.7$s.
+(b) Fast ECoT (Async) at $t=21.4$s.
+Figure 7: Qualitative examples in LIBERO experiments. Fast ECoT (Async) reaches comparable task states much quicker.
+V-B Real-world Experiments
 Setup.
@@
 We design six manipulation tasks representative of common household scenarios.
-The evaluation includes both in-distribution and out-of-distribution tasks featuring unseen objects and instructions, detailed in Sec.
-B.2.
-We collect 50 expert demonstrations via Droid teleoperation pipeline [35] to fine-tune ECoT models.
+The evaluation includes both in-distribution and out-of-distribution tasks featuring unseen objects and instructions.We collect 50 expert demonstrations via Droid teleoperation pipeline [44] to fine-tune ECoT models.
 We follow the same training data generation process, training recipe, and baselines used in Sec.
-5.1.
+V-A.
 Method ID SR (%) $\uparrow$ OOD Objects SR (%) $\uparrow$ OOD Instruction SR (%) $\uparrow$ Average SR (%) $\uparrow$ Latency per Step (ms) $\downarrow$
+OpenVLA 83.3 68.3 42.7 64.7 196 $\pm$ 32
 ECoT 78.3 73.3 40.0 64.0 5556 $\pm$ 384
@@
 ECoT (Quant) 75.0 63.3 48.3 62.2 2437 $\pm$ 171
-Fast ECoT 76.6 66.7 46.7 63.3 2479 $\pm$ 520
-Fast ECoT (Async) 83.3 76.7 50.0 70.0 716 $\pm$ 529
-Table 2: Real-world experimental results on selected household tasks.
-SR = Success Rate.
-ID = In distribution.
-OOD = Out of distribution.
-All results are averaged over 180 trials and run on a single NVIDIA RTX 4090 GPU.
-Best results are in bold.
+Fast ECoT 81.6 73.3 50.0 68.3 2479 $\pm$ 520
+Fast ECoT (Async) 78.3 75.0 42.7 65.3 716 $\pm$ 529
+TABLE II: Real-world experimental results on selected household tasks. SR = Success Rate. ID = In distribution. OOD = Out of distribution.
 Results.
 Tab.
-2 summarises the performance of our method against baseline approaches on real-world tasks.
-Fast ECoT and Fast ECoT (Async) both demonstrate substantial latency reductions while preserving or even improving task success rates.
-Specifically, Fast ECoT (Async) delivers the best performance, with an average success rate of 70% across in-distribution (ID), out-of-distribution (OOD) objects, and OOD instruction tasks, alongside the lowest latency of $716\pm 529$ ms per step—a 7.5x speedup over the original ECoT baseline ($5556\pm 384$ ms).
-This result reinforces the benefits of asynchronous parallel reasoning in real-world deployment scenarios, where low-variance and rapid inference are critical.
-5.4 Action Faithfulness in ECoT Reasoning
-Figure 6: Action faithfulness (AF) on LIBERO tasks. All graphs are plotted with mean and standard deviation shading across 1000 steps.
+II summarises the performance of our method against baseline approaches on real-world manipulation tasks.
+Fast ECoT achieves the highest overall success rate (68.3%), with strong improvements in OOD instruction tasks (50.0%, +10 points over OpenVLA).
+Meanwhile, the Async version of Fast ECoT yields the best trade-off between performance and efficiency: although its average success rate is slightly lower (65.3%), it delivers the lowest latency of $716\pm 529$ ms per step—a $7.7\times$ speedup compared to the original ECoT baseline ($5556\pm 384$ ms).
+This highlights the benefit of asynchronous parallel reasoning in real-world deployment, where rapid, low-variance inference is critical.
+In contrast, prior ECoT variants either sacrifice task performance (ECoT 5-step) or still suffer from high inference latency (ECoT and ECoT-Async), making them less practical for on-robot use.
+Consistent with the simulation examples, Fig.
+8 shows that both methods produce coherent, grounded reasoning; again, the asynchronous variant reaches comparable task states much quicker.
+(a) ECoT at $t=84.8$s.
+(b) Fast ECoT (Async) at $t=18.7$s.
+Figure 8: Real-world qualitative examples. Fast ECoT (Async) reaches comparable task states much quicker.
+V-C AutoEval Real-world Experiments
+Setup.
+We use AutoEval [45], an online platform for standardised real-robot policy evaluation.
+We choose AutoEval both for reproducible, third-party benchmarking and because its tasks/environments are drawn from BridgeData V2 [5]—the same distribution on which ECoT [4] was trained.
+This lets us run the vanilla ECoT without any finetuning and directly apply our accelerations (Fast ECoT and Fast ECoT (Async)) as drop-in replacements.
+Concretely, we evaluate four tabletop tasks in two accessible environment (Drawer, Sink) with 10 trials per task.
+Method Drawer SR (%) $\uparrow$ Sink SR (%) $\uparrow$ Latency (ms) $\downarrow$
+ECoT 30 30 4030 $\pm$ 270
+Fast ECoT 30 35 2105 $\pm$ 324
+Fast ECoT (Async) 20 25 790 $\pm$ 331
+TABLE III: AutoEval results on BridgeData V2 tabletop tasks. SR = Success Rate.
+Results.
+Fast ECoT and Fast ECoT (Async) achieve performance comparable to ECoT under AutoEval (Tab.
+III): Fast ECoT matches ECoT on Drawer (30%) and slightly exceeds it on Sink (35%), while the asynchronous variant is modestly lower (20–25%) but offers a favourable accuracy–efficiency trade-off.
+Note that AutoEval’s environment reset policy introduces distribution shift—e.g., altered object poses/orientations, handle angles, and initial camera/scene configurations—relative to the canonical starts used in the original evaluation.
+These OOD initialisations make perception and short-horizon control more volatile, reducing absolute SRs across all methods (including ECoT) even though their relative behaviour remains similar.
+V-D Updating Frequency of Across-level Reasoning
+Reusing past reasoning steps yields a temporal smoothing effect that dampens noisy fluctuations in planning and stabilises control.
+We perform a controlled ablation on LIBERO-Object (see Tab.
+IV) to examine performance differences resulting from varying the update frequency of high-level and low-level reasoning steps.
+When high-level reasoning is refreshed less frequently (every 5 frames), while low-level reasoning remains updated at every step, success rates increase from $77\%$ (baseline with updates at every step) to $79\%$.
+In Fast ECoT, each reasoning step similarly leverages current visual observations combined with cached reasoning from previous steps, achieving an even higher success rate of $83\%$.
+However, excessively infrequent updates reduce performance.
+Updating both high-level and low-level reasoning every 5 frames lowers success to $70\%$, and completely infrequent updates ($\infty$$\infty$) dramatically drop success to $35\%$.
+Common failures in such scenarios involve delayed task transitions—for instance, after grasping an object, the robot may stall indefinitely rather than moving promptly to the placement step.
+This underscores the critical balance between beneficial temporal smoothing and necessary action responsiveness.
+High-level step Low-level step Success (%) $\uparrow$
+1 1 77
+5 1 79
+5 2 78
+5 3 75
+5 5 70
+$\infty$ $\infty$ 35
+Fast ECoT Fast ECoT 83
+TABLE IV: Ablation on update frequency.
+Our findings (see Tab.
+IV) align with the results in SmolVLA [37] and Tab.
+2 of the original ECoT work, where the reuse of historical reasoning similarly improved task performance.
+Thus, moderate reuse of recent reasoning can maintain or even enhance robot performance.
+V-E Action Faithfulness in ECoT Reasoning
+Figure 9: Action faithfulness (AF) on LIBERO tasks. All graphs are plotted with mean and standard deviation shading across 1000 timesteps.
 To validate that the generated chain-of-thought (CoT) not only improves performance but also faithfully explains the model’s decision-making process, we build on the faithfulness criteria for language models [46] and introduce a novel quantitative metric - action faithfulness (AF) - to measure the faithfulness of the CoT reasoning for robotic tasks.
-fFormally, given a complete CoT consisting of $N$ intermediate reasoning steps that culminate in a final robot action $A$, we enforce the model to directly predict an action $A_{i}$ after generating only the first $i$ reasoning steps, where $i=0,1,\dots,N$.
+Formally, given a complete CoT consisting of $N$ intermediate reasoning steps that culminate in a final robot action $A$, we enforce the model to directly predict an action $A_{i}$ after generating only the first $i$ reasoning steps, where $i=0,1,\dots,N$.
 We then compute the L1 distance between $A_{i}$ and $A$ as the faithfulness score $AF_{i}=\|A_{i}-A\|_{1}$.
@@
 We plot action faithfulness scores of Fast ECoT and ECoT for all reasoning steps $\{AF_{i}\}$ in Fig.
-6.
+9.
 Fast ECoT preserves the faithfulness of the base ECoT model during parallel reasoning generation for most reasoning steps.
@@
 Faithfulness is notably low when no reasoning steps are generated (i.e., at “Task”), likely due to fine-tuning from the OpenVLA [1] checkpoint, which did not require reasoning steps.
-We provide more CoT reasoning examples in Sec.
-C.1.
-6 Conclusion
+VI CONCLUSION
 We present Fast ECoT, an inference-time acceleration method for Embodied Chain-of-Thought (ECoT) reasoning in VLA models.
@@
 Across simulations and real-world tests, it reduces inference latency by up to 7.5× while maintaining task performance and interpretability, advancing ECoT toward real-time deployment.
-7 Limitations
-While our work demonstrated that ECoT reasoning can be significantly accelerated through the reuse of previously computed thoughts, there remain several limitations.
-First, we still update all reasoning steps at once without identifying and prioritising the more vital steps for frequent updates.
-This approach can lead to excessive computation, and more sophisticated methods could dynamically determine which reasoning steps require updating based on their relative importance to task success.
-Second, the effectiveness of thoughts reuse inherently relies on the assumption that at least part of the reasoning remains stable over a short time window.
-In scenarios where reasoning steps frequently change between consecutive time steps, this stability assumption breaks down, potentially degrading the performance gains offered by our method.
-Third, our current approach assumes a fixed structure in reasoning chains, thus limiting its applicability to tasks or environments requiring highly dynamic or adaptive reasoning patterns.
-Extending thoughts reuse to accommodate more diverse and adaptive reasoning structures remains an open challenge.
-Addressing these limitations represents promising avenues for future research to further enhance the efficiency and generality of embodied reasoning in robotic control.
-Acknowledgments
-If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments.
-All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
-Appendix A Extra Algorithm Details
-A.1 Algorithm on Embodied Chain-of-Thought
-Algorithm 2 Embodied Chain‑of‑Thought
-1:Time step $t$, Instruction $I^{t}$, Observation $O^{t}$
-2:$c^{t}\leftarrow\mathrm{encode}(I^{t},O^{t})$
-3:$R^{t}\leftarrow[\,]$
-4:for $i=1\,\mathbf{to}\,N+1$ do
-5:     Autoregressively sample $r_{i}^{t}\sim P\bigl{(}r_{i}^{t}\mid c^{t},R^{t}\bigr{)}$
-6:     Add $r_{i}^{t}$ to $R^{t}$
-7:end for
-[…]
-9:return $R^{t},A^{t}$
-A.2 Algorithm on Parallel Embodied Chain-of-Thought
-Algorithm 3 Parallel Embodied Chain‑of‑Thought
-1:Time step $t$, Instruction $I^{t}$, Observation $O^{t}$, Last reasoning steps $R^{t-1}$
-2:$c^{t}\leftarrow\mathrm{encode}(I^{t},O^{t})$
-3:$R^{t}\leftarrow[\,]$
-4:for $i=1\,\mathbf{to}\,N+1$ concurrently do
-5:     Autoregressively sample $r_{i}^{t}\sim P\bigl{(}r_{i}^{t}\mid c^{t},R^{t-1}[:i]\bigr{)}$
-6:end for
-7:Synchronize
-[…]
-9:return $R^{t},A^{t}$
-Appendix B Extra Experimental Setup
-B.1 Finetuning Hyperparameters
-Hyperparameter Value
-Base Model ecot-openvla-7b-oxe
-Number of GPUs 4
-Batch Size 1
-Gradient Accumulation Steps 1
-Learning Rate 5e-4
-Learning Rate Scheduler constant LR
-Optimizer AdamW
-AdamW Betas (0.9, 0.999)
-AdamW Epsilon 1e-8
-Training Steps 200,000
-Image Augmentation
-Reasoning Dropout 0.0
-LoRA Rank 32
-LoRA Alpha 16
-LoRA Dropout 0.0
-Table 3: Fine-tuning hyperparameters used for fine-tuning EcoT model.
-B.2 Real-World Experimental Setup
-Out-of-distribution Instructions.
-We use GPT-4o [37] to rephrase instructions from the training set.
-These rephrasings modify both the verbs and nouns to create linguistically novel yet semantically equivalent instructions, allowing us to test generalization to unseen language.
-Original Instruction OOD Rephrased Instruction
-Pick up the gray spanner and put it in the basket Grab the silver wrench and drop it into the bin
-Pick up the blue hammer and put it in the basket Take the cobalt mallet and place it in the container
-Pick up the blue squash ball and put it in the basket Retrieve the navy rubber ball and set it in the basket
-Pick up the white baseball and put it in the basket Grab the pale softball and store it in the bucket
-Pick up the banana and put it on the plate Take the yellow fruit and lay it on the dish
-Pick up the lemon and put it in the cup Seize the citrus and drop it into the mug
-Table 4: Original instructions and their corresponding out-of-distribution (OOD) rephrasings generated by GPT-4o.
-In-distribution Objects.
-Figure 7: Qualitative examples of in-distribution (ID) objects used for evaluation.
-Out-of-distribution objects.
-Figure 8: Qualitative examples of out-of-distribution (OOD) objects used for evaluation.
-Appendix C Extra Experimental Results
-C.1 CoT Reasoning Examples
-Figure 9: Qualitative examples of ECOT reasoning in real-world experiments.
-Figure 10: Qualitative examples of ECOT reasoning in simulation.
-As the VLMs used in both the data generation pipeline and the underlying VLA model are pretrained on real-world images, their performance in simulated environments is less stable.
-This can lead to inaccurate bounding boxes, undetected objects, and unusual or incorrect object labels.